Systems Staff Meeting Notes
January 25, 2008
January 25, 2008
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Stephen WilliamsAnnouncements
Ongoing Systems Updates
- The Hogtown Medieval Festival is this weekend.
- Library Council - discussion of new reorganization chart. Judy will be discussing it with all staff at today's Town Meeting. We also discussed how there will be another budget reduction as of July 1st.
- Judy will be attending our next staff meeting.
- Software orders from SLS - very confusing. Will is trying to sort them out.
- The old Web server has been rebuilt. It needs the operating system reinstalled. We will use it as a VM server.
- Wireless in MSL - Amy worked hard to get the laptops functional, but now they may be sent back to GovConnection. The battery charger they sold us is not compatible with the laptop batteries. If they cannot find a solution to the problem, we will send them back and/or get a refund.
- Systems training is going well. Group 3 is scheduled for next week, and then we will change topics and start again.
Software Projects Status
- McAfee test server looks good. Still testing, but workstations are being updated.
- SAN - still awaiting licenses. Vendor meeting schedule for Feb. 12. We've had some issues with Replistor over the last couple of days - problems mirroring to the backup server. Could be a bandwidth issue. CNS upgraded the feeds to the server at MSL to 1 Gig. DLC reports a slowness issue with the SAN - will have them log the time the problems occur.
- SMS upgrade - database issue with connection to the SQL server.
- SAN move to CSE is probably not going to happen - the charge would be too hefty - over $500 per month. Bill wants us to scout out MSL for a different space - maybe the Xerox office.
- Central Exchange server - nothing new. Meetings are scheduled for the near future.
- Campus AD print demo - scheduled for Feb. 6. The migration of public computers to Campus AD has been postponed beyond spring break - to Summer A. Working on a workaround for the 14-digit logon as well as an invisible patron tracking program. Will consider notifying users of tracking.
- Consolidated logging - Logan met with John Sawyer and got some ideas. This will be on hold for a few weeks.
- Discussion about synching My Docs to home folder on the SAN.
DLC Greenstone Issues
- Patron Logging - Stephen is working on log on/off scripts. It will be invisible.
- Stephen gave a great presentation on blogs and wikis at the cross-divisional meeting.
- Grover is ongoing.
- Tom Bielicke - latest version of DLC importer released for testing. Erich is anxious for it - should be a quick test.
- Greenstone - optimization phase - Mark is working on memory and speed.
- SL2K3DLC1 - working on permissions for Erich. He wants users to have modify, but not delete. Winston is testing - unsure if that is possible.
- Wiki update - Jason Fleming really wants to get going on this. Stephen is trying to get authentication working.
- SQL server - Winston is adjusting the permissions so that Domain Admins can't create databases. Creating databases or adjusting the size on the SQL server must go through Winston.
- SMS - WSUS reinstalled. Updates approved.
- FTP - Logan is working on it.
- SCSI card installed.
- Added 2TB drive. Another drive has not been mounted.
Systems Security Management
- Server-side includes to be turned on next week after the weekly backups run. FrontPage extensions will need to be turned off. Web Content Providers folks were notified. After that message went out, there were requests to modify the members of that distribution list.
- Meeting held.